write (2-i)(4-3i)^2/3+2i in the form a+bi where a,b,er.


1
Expert's answer
2022-05-27T10:39:23-0400

"\cfrac{(2-i)(4-3i)^2}{3+2i} = \cfrac{(2-i)(16-24i-9)}{3+2i} =\cfrac{(2-i)(7-24i)}{3+2i} = \\ 
=\cfrac{14-48i-7i-24}{3+2i}
=\cfrac{-10-55i}{3+2i} \\
=\cfrac{-10\cdot3 -55\cdot2}{9+4}+\cfrac{-10\cdot2-55\cdot3}{9+4}i= -\cfrac{140}{13}-\cfrac{145}{13}i"